OpenStudy (anonymous):

Andrew thinks of 4 consecutive numbers. The sum of the first 3 numbers is 100 more than the fourth number. What is the sum of these 4 consecutive numbers?

OpenStudy (anonymous):

If x is one consecutive number than tell me the next number consecutive to it..

OpenStudy (anonymous):

x+1

OpenStudy (anonymous):

And others will be??

OpenStudy (anonymous):

x + (x+1) + (x+2) = (x+3)+100

OpenStudy (anonymous):

now solve for x

OpenStudy (anonymous):

Yes now solve for x here @Genuine

OpenStudy (anonymous):

Let him do @sauravshakya

OpenStudy (anonymous):

and find: x + (x+1) + (x+2)+(x+3) which will be your answer

OpenStudy (anonymous):

Oh i get it now, I can solve it.

OpenStudy (anonymous):

What is x value you got tell us here after solving.. Take your time..

OpenStudy (anonymous):

x=50

OpenStudy (anonymous):

Yes now find the sum of all the numbers and then tell us..

OpenStudy (anonymous):

50+51+52+53=206

OpenStudy (anonymous):

Well Done..

OpenStudy (anonymous):

A medal for your work..

OpenStudy (anonymous):

Thankyou very much!

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

\[\large \boxed {\color{green}{\textbf{Well Done.. @Genuine}}}\]